Suppose that you are planning a two-group experiment. How high would your t value have to be for half of the variance of your dependent variable to be accounted for

a. If you use 25 participants per group?

b. If you use 81 participants per group?

Please show ALL work, including the formulas used.
Answer #1

a)here number of independent variables k=1

n=25

variance explained =R2 =0.5

F=((n-k-1)*R2/(k*(1-R2))=(25-1-1)*0.5/(1*(1-0.5))= 23.000

therefore value of t =sqrt(F) =sqrt(23)=4.796

b)n=81

F=((n-k-1)*R2/(k*(1-R2))= 79.000

therefore value of t =sqrt(F) =sqrt(79)=8.888